Find two numbers whose sum is -2 and whose difference is 35 by setting up a system of linear equations and solving it. Remember to appropriately define and variables you use.
Answer by stanbon(75887)

Find two numbers whose sum is -2 and whose difference is 35 by setting up a system of linear equations and solving it. -----
Equations:
x + y = -2
x - y = 35
-----
Add and solve for "x"::
2x = 33
x = 16.5
-----
Solve for "Y":
x + y = -2
16.5 + y = -2
-----
y = -18.5
-----------------------
Cheers,
Stan H.